# Analytic approach to boundary integrability with application to mixed-flux $AdS_{3}\times S^{3}$

### v1

@@ II Mixed-flux $AdS_{3}\times S^{3}$
Here $q_{\text{\tiny R}},q_{\text{\tiny NS}}\in\mathbb{R}$ denote the amount of RR and NSNS flux respectively, with $q_{\text{\tiny R}}^{2}+q_{\text{\tiny NS}}^{2}=1$.
@@
Both ${\cal L}(\mathsf{x})$ and ${\cal L}^{g_{\text{\tiny B}}}(\mathsf{x})$ have the same pole divisor: two simple poles at $\mathsf{x}_{\pm}=\pm 1$. In addition, they each have a single, but different, simple zero. In the above coset-gauge, ${\cal L}(\mathsf{x}_{0})=0$ and ${\cal L}^{g_{\text{\tiny B}}}(\mathsf{x}_{0}^{g})=0$ with
@@
Notice that when keeping the coset-gauge for $\tilde{g}_{\text{\tiny B}}$ general, ${\cal L}(\mathsf{x})$ would not have a zero, while ${\cal L}^{g}(\mathsf{x})$ remains zero at $\mathsf{x}=\mathsf{x}_{0}^{g}$ (which reduces to $\mathsf{x}=\infty$ in the pure RR case). This corresponds to the standard set-up for coset sigma-models where one expands the Lax representative ${\cal L}^{g}(\mathsf{x})$ to obtain local charges. The possibility of having an additional zero when using $\tilde{g}_{\text{\tiny B}}=\mathbb{1}$ is however special to the $AdS_{3}\times S^{3}$ (permutation-)coset case and is important for its integrable boundary structure.
Given (6), there are then four possible reflections $\rho_{i}(\mathsf{x})$, depending on whether the poles and zeros are fixed or permuted. The behaviour of the zeros further determines the gauges for the Lax connections: when $\rho(\mathsf{x}_{0})=\mathsf{x}_{0}$ we must take $U=\mathbb{1}$, while for $\rho(\mathsf{x}_{0})=\mathsf{x}_{0}^{g}$ we must take $U=g_{\text{\tiny B}}$. [^5] Note that $\rho(\mathsf{x}_{0}^{g})=\mathsf{x}_{0}^{g}$ with both Lax representatives in the gauge $U=g_{\text{\tiny B}}$ is equivalent to $\rho(\mathsf{x}_{0})=\mathsf{x}_{0},U=\mathbb{1}$ up to redefining $K_{\sigma_{\star}}$. In table 1, we summarise these possibilities, the corresponding solutions of (7) and boundary conditions from (4) for trivial $K_{\sigma_{\star}}$ matrices.
@@ 4 $\rho(\mathsf{x}_{0})=\mathsf{x}^{g}_{0}$ $\rho(\mathsf{x}_{\pm})=\mathsf{x}_{\pm}$
Specifically, one finds that there is no solution $\rho_{4}(\mathsf{x})$ fixing the poles and permuting the zeros. Yet both poles and zeros are fixed by $\rho_{3}(\mathsf{x})=\mathsf{x}^{-1}$ when $q_{\text{\tiny R}}=0$. However, in this case the Lax connection does not depend on $\mathsf{x}$ and for trivial $K_{\sigma_{\star}}$ one has $T_{\text{\tiny b}}(\mathsf{x})=\mathbb{1}$. The boundary conditions (4) would be trivially satisfied and thus can not provide solutions of the sigma-model boundary equations of motion (LABEL:eq:sm-bcs) (see also [11]).
The interesting cases arise when poles are permuted. Firstly, we have
@@
which fixes $\mathsf{x}_{0}$. This reduces to the standard reflection $\rho_{1}(\mathsf{x})=-\mathsf{x}$ in the pure RR case ($q_{\text{\tiny NS}}=0$). Mapping residues of the minimal boundary conditions ${\cal L}_{\tau}(\mathsf{x})\overset{\text{\tiny b}}{=}{\cal L}_{\tau}(\rho_{1}(\mathsf{x}))$ then in general gives
@@
where $\mu_{\pm}$ has cardinality $m_{\pm}$. For the automorphism given in (39), one finds e.g. $m_{+}=4$ and then this corresponds to a D-brane with $p=3$ and ${\cal F}=0$. Nontrivial worldvolume fluxes are typically introduced by dynamical and $\mathsf{x}$-dependent $K_{\sigma_{\star}}$-matrices, as happens for several interesting D-branes in $AdS_{5}\times S^{5}$ and $AdS_{4}\times\mathbb{CP}^{3}$ [30, *Linardopoulos:2022wol, *Demjaha:2025axy, 33]. In the future, it would be interesting to compare with the $\mathsf{x}$-dependent quantum $K_{\sigma_{\star}}$-matrices of [34, *Bielli:2024xuv, *Bielli:2025abu] known for pure RR flux. For generic mixed-flux with (10), an interesting question is also whether compatibility with (LABEL:eq:sm-bcs) requires dynamical reflection data $K_{\sigma_{\star}}$.
The second option permutes both poles and zeros
@@
implying the constraint $(W\mathrm{Ad}_{g_{\text{\tiny B}}})^{2}\overset{\text{\tiny b}}{=}\mathbb{1}$. Interestingly, for $q_{\text{\tiny R}}=0$, where ${\cal L}(\mathsf{x})$ and ${\cal L}^{g_{\text{\tiny B}}}(\mathsf{x})$ become $\mathsf{x}$-independent, the boundary conditions enhance to $\mathcal{J}^{\text{\tiny B}}_{+}=-W\left[\mathcal{R}_{-}^{\text{\tiny B}}\right]$ without further constraints. These are precisely the conformal boundary conditions of WZW models which correspond to D-branes wrapping twisted conjugacy classes in $SU(1,1)\times SU(2)$. For $W=\mathbb{1}$ this yields unstable $dS_{2}\times S^{2}$ branes, which we will exclude in the following, while including the nontrivial outer automorphism (40) with $K_{\sigma_{\star}}=w_{\text{\tiny o}}$ yields stable $AdS_{2}\times S^{2}$ branes [25]. In terms of the fields $X^{\mu}=\{\psi,\omega,\tau;\alpha,\beta,\gamma\}$ given in (36), the latter boundary conditions read
@@
Before doing so, we remark that interestingly $\rho_{1}(\mathsf{x})$ at $q_{\text{\tiny NS}}=0$ can equally be interpreted as demanding the $\sigma\rightarrow 2\pi-\sigma$ parity-invariance of ${\cal L}(\mathsf{x})$ after reflection, while $\rho_{2}(\mathsf{x})$ combines requiring invariance under both parity and group element inversion $g_{\text{\tiny B}}\rightarrow g_{\text{\tiny B}}^{-1}$. It would be interesting to relate this to crossing properties of [15, *Lloyd:2014bsa, *OhlssonSax:2023qrk, *Frolov:2025uwz, *Frolov:2025tda].
@@
with $T_{k}$ the $\mathfrak{su}(1,1)\oplus\mathfrak{su}(2)$ generators given in (27) and $u_{k}$ depending on $\tau$ only through the target-space fields $X^{\mu}=\{\psi,\omega,t;\alpha,\beta,\gamma\}$. Projecting the resulting system (4) onto the generators yields six coupled first-order PDEs in the worldsheet variables $(\tau,\sigma)$. Using (15), we rewrite them in terms of $\partial_{\sigma}\{\psi,\alpha\}$ and $\partial_{\tau}\{t,\omega,\beta,\gamma\}$ only. Imposing that $u_{k}$ does not depend on $\partial_{\sigma}\{\psi,\alpha\}$ (i.e. they can depend on $\psi,\alpha$ but not their derivatives), gives a set of algebraic conditions solved by
@@
When the denominator of $z(\mathsf{x})$ vanishes, all $u_{k}=0$ and $K_{\sigma_{\star}}=w_{\text{\tiny o}}$. This occurs at $\mathsf{x}=\mathsf{x}_{0}$ and at the WZW point. Moreover also on the loci (17) we get $K_{\sigma_{\star}}=w_{\text{\tiny o}}$. This is thus entirely consistent with the previous subsection. Away from these special points, we thus find that the brane embedding can remain that of a twisted conjugacy class as in boundary WZW models, while $K_{\sigma_{\star}}(\mathsf{x})$ becomes nontrivial showing that the mixed-flux deformation is encoded entirely in the reflection data.
@@ Acknowledgements.
We thank D. Bielli, C. Kristjansen, S. Lacroix, G. Linardopoulos, B. Nairz, D. Polvara, A. Retore, K. Zarembo for useful discussions, and A. Retore, K. Zarembo and especially S. Lacroix for comments on the draft. This work is partly based on the Master thesis of JCG prepared at ETH Zürich and we thank N. Beisert for acting as co-advisor. JCG’s research is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - Projektnummer 417533893/GRK2575 “Rethinking Quantum Field Theory”. SD is supported by the Swiss National Science Foundation through the SPF fellowship TMPFP2_224600\_224600 and the NCCR SwissMAP. SD also thanks the organisers and participants of the workshop Higher-$d$ integrability (Favignana, 2025) where part of this work was presented.
@@ Appendix B Mixed-flux $AdS_{3}\times S^{3}$ bulk sigma-model
In this appendix, we collect the ingredients for the boundary integrability analysis of the mixed NSNS/RR flux string on $AdS_{3}\times S^{3}\times M_{4}$, with $M_{4}$ a four-dimensional compact manifold required to complete the type IIB supergravity background (see [12, *Seibold:2024qkh] for reviews). Since the integrable structure relevant for our purposes is encoded in the $AdS_{3}\times S^{3}$ sector, we will restrict attention to its Green-Schwarz formulation as a $\mathbb{Z}_{4}$-graded supercoset sigma-model [14] (see also [37, *Hoare:2014oua]), while keeping $M_{4}$ implicit throughout.

### v2

@@ II Mixed-flux $AdS_{3}\times S^{3}$
Here $q_{\text{\tiny R}},q_{\text{\tiny NS}}\in\mathbb{R}$ denote the amount of RR and NSNS flux of the full theory respectively, related as $q_{\text{\tiny R}}^{2}+q_{\text{\tiny NS}}^{2}=1$.
@@
Both ${\cal L}(\mathsf{x})$ and ${\cal L}^{g_{\text{\tiny B}}}(\mathsf{x})$ have the same pole divisor: two simple poles at $\mathsf{x}_{\pm}=\pm 1$. In addition, they each have a single, but different, simple zero. In the above coset-gauge, ${\cal L}(\mathsf{x}_{0})=0$ and ${\cal L}^{g_{\text{\tiny B}}}(\mathsf{x}_{0}^{g})=0$ with [^7] When keeping the coset-gauge for $\tilde{g}_{\text{\tiny B}}$ general, ${\cal L}(\mathsf{x})$ would have no zero while ${\cal L}^{g}(\mathsf{x})$ remains zero at $\mathsf{x}=\mathsf{x}_{0}^{g}$ (which in the pure RR case reduces to $\mathsf{x}=\infty$). This corresponds to the standard set-up for coset models where one expands ${\cal L}^{g}(\mathsf{x})$ to obtain local charges. The possibility of having an additional zero when using $\tilde{g}_{\text{\tiny B}}=\mathbb{1}$ is however special to $AdS_{3}\times S^{3}$ (permutation-)cosets.
@@
Given (6), there are then four possible reflections $\rho_{i}(\mathsf{x})$, depending on whether the poles and zeros are fixed or permuted. The behaviour of the zeros further determines the necessary gauges for the Lax connections: when $\rho(\mathsf{x}_{0})=\mathsf{x}_{0}$ we must take $U=\mathbb{1}$, while for $\rho(\mathsf{x}_{0})=\mathsf{x}_{0}^{g}$ we must take $U=g_{\text{\tiny B}}$. [^8] Note that $\rho(\mathsf{x}_{0}^{g})=\mathsf{x}_{0}^{g}$ with both Lax representatives in the gauge $U=g_{\text{\tiny B}}$ is equivalent to $\rho(\mathsf{x}_{0})=\mathsf{x}_{0},U=\mathbb{1}$ up to redefining $K_{\sigma_{\star}}$. In table 1, we summarise the possibilities, the solutions of (7) and the “minimal” boundary conditions from (4), i.e. for trivial $K_{\sigma_{\star}}$ matrices.
@@ 4 $\rho(\mathsf{x}_{0})=\mathsf{x}^{g}_{0}$ $\rho(\mathsf{x}_{\pm})=\mathsf{x}_{\pm}$
Specifically, one finds that there is no solution $\rho_{4}(\mathsf{x})$ fixing the poles and permuting the zeros. Yet both poles and zeros are fixed by $\rho_{3}(\mathsf{x})=\mathsf{x}^{-1}$ when $q_{\text{\tiny R}}=0$. However, in this case the Lax connection does not depend on $\mathsf{x}$ and for trivial $K_{\sigma_{\star}}$ one has $T_{\text{\tiny b}}(\mathsf{x})=\mathbb{1}$ and trivial boundary conditions (4) that can not provide solutions of the sigma-model boundary equations of motion (LABEL:eq:sm-bcs).
The interesting cases arise when poles are permuted: Branch 1. Firstly, we have
@@
which fixes $\mathsf{x}_{0}$. For $q_{\text{\tiny NS}}=0$, this reduces to the standard reflection $\rho_{1}(\mathsf{x})=-\mathsf{x}$ obtained from parity-invariant gluing in pure RR cosets [7, 30, *Linardopoulos:2022wol, *Demjaha:2025axy]. Mapping residues of the conditions ${\cal L}_{\tau}(\mathsf{x})\overset{\text{\tiny b}}{=}{\cal L}_{\tau}(\rho_{1}(\mathsf{x}))$ in general now gives
@@
where $\mu_{\pm}$ has cardinality $m_{\pm}$. For the automorphism given in (39), one finds e.g. $m_{+}=4$ and then this corresponds to a D-brane with $p=3$ and ${\cal F}=0$. Nontrivial worldvolume fluxes are typically introduced by dynamical and $\mathsf{x}$-dependent $K_{\sigma_{\star}}$-matrices, as happens for several interesting D-branes in $AdS_{5}\times S^{5}$ and $AdS_{4}\times\mathbb{CP}^{3}$ [30, *Linardopoulos:2022wol, *Demjaha:2025axy, 33]. In the future, it would be interesting to compare with the $\mathsf{x}$-dependent quantum $K_{\sigma_{\star}}$-matrices of [34, *Bielli:2024xuv, *Bielli:2025abu] known for pure RR flux. For nontrivial mixed-flux, an interesting question is also whether compatibility between (10) and (LABEL:eq:sm-bcs) requires dynamical reflection data $K_{\sigma_{\star}}$.
Branch 2. Both poles and zeros are permuted by
@@
implying the constraint $(W\mathrm{Ad}_{g_{\text{\tiny B}}})^{2}\overset{\text{\tiny b}}{=}\mathbb{1}$. Interestingly, for $q_{\text{\tiny R}}=0$, where ${\cal L}(\mathsf{x})$ and ${\cal L}^{g_{\text{\tiny B}}}(\mathsf{x})$ become $\mathsf{x}$-independent, the boundary conditions enhance to $\mathcal{J}^{\text{\tiny B}}_{+}=-W\left[\mathcal{R}_{-}^{\text{\tiny B}}\right]$ without further constraints. These are precisely the conformal boundary conditions of WZW models which correspond to D-branes embedded as twisted conjugacy classes in $SU(1,1)\times SU(2)$. For $W=\mathbb{1}$ this yields unstable $dS_{2}\times S^{2}$ branes, which we will exclude in the following, while including the nontrivial outer automorphism (40) with $K_{\sigma_{\star}}=w_{\text{\tiny o}}$ yields stable $AdS_{2}\times S^{2}$ branes [25]. In terms of the fields $X^{\mu}=\{\psi,\omega,\tau;\alpha,\beta,\gamma\}$ given in (36), the latter boundary conditions read
@@
Before doing so, we remark that, interestingly, $\rho_{1}(\mathsf{x})$ at $q_{\text{\tiny NS}}=0$ can equally be interpreted as demanding the $\sigma\rightarrow 2\pi-\sigma$ parity-invariance of ${\cal L}(\mathsf{x})$ after reflection, while $\rho_{2}(\mathsf{x})$ combines requiring invariance under both parity and group element inversion $g_{\text{\tiny B}}\rightarrow g_{\text{\tiny B}}^{-1}$. It would be interesting to relate this to crossing properties of [15, *Lloyd:2014bsa, *OhlssonSax:2023qrk, *Frolov:2025uwz, *Frolov:2025tda].
@@
with $T_{k}$ the $\mathfrak{su}(1,1)\oplus\mathfrak{su}(2)$ generators given in (27) and $u_{k}$ depending on $\tau$ only through the target-space fields $X^{\mu}=\{\psi,\omega,t;\alpha,\beta,\gamma\}$. Projecting the resulting system (4) onto the generators yields six coupled first-order PDEs in the worldsheet variables $(\tau,\sigma)$. Using (15), we rewrite them in terms of $\partial_{\sigma}\{\psi,\alpha\}$ and $\partial_{\tau}\{t,\omega,\beta,\gamma\}$ only. Requiring that $u_{k}$ does not depend on $\partial_{\sigma}\{\psi,\alpha\}$ (i.e. $u_{k}$ depends on the fields but not their derivatives), gives a set of algebraic conditions solved by
@@
We obtain $u_{k}=0$ and hence $K_{\sigma_{\star}}=w_{\text{\tiny o}}$ on the loci (17) and whenever the denominator of $z(\mathsf{x})$ vanishes, which occurs at $\mathsf{x}=\mathsf{x}_{0}$ and at $q_{\text{\tiny R}}=0$ (the WZW point). This is therefore fully consistent with the previous subsection. Away from these special points, we find that branes wrapping twisted conjugacy classes, as in boundary WZW models, remain compatible with integrability, with the mixed-flux deformation carried entirely by the nontrivial reflection data $K_{\sigma_{\star}}(\mathsf{x})$ rather than the brane geometry.
@@ Acknowledgements.
We thank D. Bielli, C. Kristjansen, S. Lacroix, G. Linardopoulos, B. Nairz, D. Polvara, A. Retore, and K. Zarembo for useful discussions, and A. Retore, K. Zarembo and especially S. Lacroix for comments on the draft. This work is partly based on the Master thesis of JCG prepared at ETH Zürich and we thank N. Beisert for acting as co-advisor. JCG’s research is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - Projektnummer 417533893/GRK2575 “Rethinking Quantum Field Theory”. SD is supported by the Swiss National Science Foundation through the SPF fellowship TMPFP2_224600\_224600 and the NCCR SwissMAP. SD also thanks the organisers and participants of the workshop Higher-$d$ integrability (Favignana, 2025) where part of this work was presented.
@@ Appendix B Mixed-flux $AdS_{3}\times S^{3}$ bulk sigma-model
In this appendix, we collect the relevant ingredients for the boundary integrability analysis of the mixed NSNS/RR flux string on $AdS_{3}\times S^{3}\times M_{4}$, with $M_{4}$ a four-dimensional compact manifold required to complete the type IIB supergravity background (see [12, *Seibold:2024qkh] for reviews). Since the integrable structure relevant for our purposes is encoded in the $AdS_{3}\times S^{3}$ sector, we will restrict attention to its Green-Schwarz formulation as a $\mathbb{Z}_{4}$-graded supercoset sigma-model [14] (see also [37, *Hoare:2014oua]), while keeping $M_{4}$ implicit throughout.